OpenStudy (anonymous):

Hannah makes her own furniture polish using 30% vinegar and 70% olive oil. Write and solve an equation to determine how much total solution she made if she use 9 tablespoons of vinegar show your work plz help me I give medal

OpenStudy (anonymous):

Total solution must be of 30 tablespoons

OpenStudy (anonymous):

30% = 9 tbs 1% = 9/30 =3/10 70% of oil = 70 (3/10) = 7(3) = 21tbs Total no of tbs = 21+9 = 30 tbs